OpenStudy (anonymous):

Help! I can't get this question even though it appears fairly easy. The equation we're supposed to use for the speed of sound is: Speed of sound= 331+.6(T), where T is the air temperature. The temperature of the air is 11.0 °C in a room. Heat is pumped into the room to alter the air temperature. If the speed of sound changes by a factor of 1.06 due to this change in temperature, what was the final temperature in the room?

OpenStudy (anonymous):

What are units of T? oF or oC or absolute oK? v = 331 +0.6 T v' = 1.06 v = 331+ 0.6(T -11) solve for T

OpenStudy (anonymous):

The units will be Celsius.

OpenStudy (anonymous):

then v = 331 + 0.6(11) = 331+6.6= 337.6 1.06 v = 357.9 = 331 + 0.6(T) 0.6 T = 357.9 - 331 = 26.9 T = 44.8 check: v' = 331 + (0.6)(44.8) = 357.9= (1.06)(337.6) checks my original equation was wrong: 1.06 v = 331 + 0.6(T-11)

OpenStudy (anonymous):

My apologies for original error.

OpenStudy (anonymous):

Thank you so much!! You rock!!!!!